# Many Minds from One Model: Bayesian Transformers for Population Intelligence

### v1

@@ 4.3 Temporal Consistency via Noise Caching
A requirement for valid variational approximation in sequential models is maintaining a consistent stochastic mask across time steps, as established in prior work on Recurrent Neural Networks gal2016dropout. Naively sampling $\mathbf{z}$ at every token step violates this principle, resulting in a temporally incoherent model that effectively shifts its functional hypothesis mid-sentence.
To adhere to this theoretical constraint within our framework, we implement a sequence-level noise caching mechanism (see Figure 1). As detailed in the BayesianBiasWrapper, the process ensures that the stochastic perturbation defines a stable reasoning persona for the duration of the generation: (1) Initialization: Upon receiving the prompt, we sample the latent variable $\mathbf{z}$ once and cache it, effectively drawing a single sample $\theta^{(k)}$ from the approximate posterior. (2) Generation: For all subsequent autoregressive steps within the same sequence, the model applies this frozen $\mathbf{z}$ to the normalization layers.
This design ensures that the stochasticity is strictly inter-sequence rather than intra-sequence. Crucially, unlike full weight-masking methods (e.g., MC Dropout), our implementation is computationally negligible, requiring only an element-wise addition of a cached tensor without the overhead of generating large binary masks. Specifically, implementing theoretically correct (frozen) MC Dropout on a 7B model would require caching around 14GB of masks per user instance, assuming pre-expand the binary masks to FP16 to align with weight precision to avoid prohibitive on-the-fly decompression overheads. In contrast, B-Trans only caches the bias noise vectors (<1MB), reducing the memory footprint by orders of magnitude while maintaining temporal consistency.
@@ 5 Experiments
Our evaluation probes B-Trans from three complementary angles. First, we test whether sampling model instances produces functional diversity in zero-shot generation rather than surface-level randomness. Second, we study exploration in Reinforcement Learning with Verifiable Rewards (RLVR), asking whether parameter-space sampling improves search under sparse rewards. Third, we examine population aggregation in a label-free Test-Time RL (TTRL) setting, where the learning signal is induced from model outputs rather than ground-truth supervision. Across all settings, our goal is to provide a minimal proof-of-concept for population-based exploration under controlled, standard protocols, rather than an exhaustive benchmark. Accordingly, we keep training and decoding setups aligned with established baselines and focus on whether B-Trans yields consistent improvements in exploration and semantic diversity.
It is important to note that while B-Trans modifies the architecture’s normalization layers, the modified backbone remains differentiable and trainable via standard backpropagation pipelines. From the efficiency and computing concern, we follow Low-Rank Adaptation (LoRA) hu2022lora training for the RLVR and TTRL tasks. In all training settings, we inject stochasticity only during rollout for exploration, while keeping the update phase deterministic by using the mean shift only, matching standard optimization stability requirements.
@@
Figure 4: PCA visualization of response embeddings from Qwen3 and Llama-3.1 models. The query is “Output a hard question to humanity (super concise and short), independent of theme.” Standard sampling confines models to disjoint semantic regions (their respective “territories”). B-Trans significantly expands this scope, allowing the model to escape its default mode and even explore semantic regions typically associated with the opposing model (highlighted in red box).
5.1 Zero-Shot Diversity and Creativity
The premise of B-Trans is that different instance samples correspond to distinct functional behaviors. We evaluate whether B-Trans yields meaningful heterogeneity compared to high-temperature decoding.
@@ 5.1 Zero-Shot Diversity and Creativity
We study intrinsic population behavior on two benchmarks with different measurement goals. On MMLU-Pro wang2024mmlu (reasoning), we use Pass@K to test whether additional sampled individuals provide functional and valid reasoning paths: for each question we sample $K$ independent sequence-level personas and count the answer as correct if any sample is correct. Thus, improved scaling with larger $K$ indicates that the induced diversity is not mere noise but yields additional correct solutions. On INFINITY-CHAT jiang2025artificial (open-ended creativity), where there is no ground-truth correctness, we measure semantic diversity using average pairwise embedding cosine distance reimers2019sentence; song2020mpnet and visualize the embedding geometry with PCA following Artificial Hivemind (jiang2025artificial). The goal is to quantify and visualize semantic heterogeneity under the same generation setup as generating answer for MMLU-Pro. For both tasks, we contrast B-Trans against two distinct base models, Qwen3 yang2025qwen3 and Llama-3.1 grattafiori2024llama, using high-temperature sampling ($T=1$). This setup allows us to observe whether the models can escape their default semantic modes through B-Trans versus Temperature stochasticity.
@@ Results.
On MMLU-Pro (Figure 2), while B-Trans matches the baseline at $K=1$, it demonstrates significantly stronger scaling behavior as the sample budget increases. This trend confirms that the diversity introduced by B-Trans is semantically meaningful: rather than generating random noise (which would yield diminishing returns), the implicit population successfully explores valid reasoning trajectories that are inaccessible to the deterministic model. Complementing this, the embedding distance analysis on INFINITY-CHAT (Figure 3) further quantifies this behavior, confirming superior output heterogeneity.
Qualitatively, Figure 4 visualizes the semantic landscape. Under standard sampling, Qwen3 and Llama-3.1 occupy disjoint, narrow regions in the semantic space, reflecting their distinct pre-trained biases. While high temperature slightly expands these clusters, it fails to alter their fundamental topology. In contrast, B-Trans generates a significantly broader distribution, forming distinct clusters that explore diverse reasoning paths.
Analysis.
The visualization in Figure 4 reveals a critical insight into the nature of the generated diversity. Standard temperature sampling is confined to local perturbations; it cannot push a model out of its inherent “personality,” leaving Qwen3 and Llama-3.1 isolated in their respective semantic territories. Conversely, B-Trans enables the models to transcend these native constraints. The visualization shows that B-Trans samples not only cover a wider area but also begin to bridge the gap between the two models, with samples from one model occasionally reaching the semantic territory typically occupied by the other. This suggests that perturbing normalization biases can shift the model’s generation distribution in embedding space beyond local temperature-based perturbations. Meanwhile, the overlap indicates that B-Trans can occasionally reach semantic regions that are atypical for the original base model under the same decoding budget, thus overlapping with the reasoning styles of different model families, rather than merely adding noise to a single persona.
@@ 5.2 Exploration in Reinforcement Learning with Verifiable Rewards (RLVR)
Table 1: Performance on RLVR tasks using GRPO. B-Trans consistently outperforms the baseline, particularly on smaller models and harder datasets.
In Reinforcement Learning, the critical bottleneck is often the exploration-exploitation trade-off. Standard LLM policies rely on action-space noise (sampling tokens) to explore. We hypothesize that parameter-space noise (sampling models) provides a more consistent and deeper form of exploration. By solving mathematical problems where the solution space is sparse, we test the competency of the implicit B-Trans population. If B-Trans achieves higher rewards under Group Relative Policy Optimization (GRPO) guo2025deepseek, it indicates that the diverse rollout candidates are not just different, but are hitting the correct solution regions more frequently than the baseline.
Setting.
To demonstrate the impact of B-Trans on exploration and exploitation trade-offs, we utilize the standard RLVR setting, following the SimpleRL-Zoo framework  [^2] We follow the default configuration as-is; the only change is apply BayesianBiasWrapper during the rollout, as we do not introduce any additional modifications beyond B-Trans.. To save computing, we use LoRA instead of full-parameter fine-tuning (less than 1% trainable parameters), and we train on 400 examples sampled from the same distribution of SimpleRL-Zoo dataset via difficulty-stratified sampling.
In GRPO, a group of outputs is sampled for each input, and the policy is updated based on their relative rewards. We test on four standard mathematical reasoning datasets, i.e., GSM8K lewkowycz2022solving, MATH-500 lightman2023lets, and Minerva Math lewkowycz2022solving, where the ground truth is verifiable.
Results.
We compare GRPO training with standard decoding against GRPO training where B-Trans instance sampling is enabled during rollout, under the same protocol. As shown in Table 1, B-Trans achieves consistent gains across all four benchmarks.
GRPO relies on informative relative comparisons within each sampled group. When the rollout group is homogeneous, relative advantages become less informative and updates can stagnate. By producing group members from distinct parameter-space instances, B-Trans increases the chance that at least one candidate reaches a higher-reward region, strengthening the learning signal available to GRPO under the same group size.
5.3 Test-Time RL without Labels
This experiment represents the most challenging scenario: learning without ground-truth labels. We employ Test-Time RL (TTRL), a setting where the model must self-improve by generating candidate solutions and filtering them via consistency checks. Specifically, we follow the established TTRL protocols zuo2025ttrl; yu2025restrain to utilize majority voting as the reward proxy. It is worth noting that the majority voting inherently rewards consensus, seemingly contrasting with B-Trans’s philosophy of encouraging diverse exploration (“The truth is often with the few”). However, our contribution focuses on the quality of the exploration rather than the selection mechanism. We argue that a superior learning signal should emerge from diverse independent reasoning paths rather than simple outcome consensus from mode collapse. Under the GRPO framework, standard models often generate homogeneous samples, providing redundant gradient information even when the reward is positive. Conversely, B-Trans injects parameter-space diversity, ensuring that the high-reward trajectories (those aligning with the majority) exhibit varied reasoning paths. This semantic diversity maximizes the information gain per update step, allowing the model to explore diverse reasoning patterns rather than overfitting to a single, narrow solution path.
Setting.
We investigate the efficacy of B-Trans in scenarios characterized by sparse signals and challenging search spaces using the AIME24 dataset. Unlike standard RLHF, TTRL adapts the model using only unlabeled test prompts. To synthesize a reward signal without ground truth, we utilize an outcome-based Majority Voting mechanism. For each input, we sample $N$ candidate responses and extract the final numerical answer. A consensus label is determined solely by the most frequent final answer. Trajectories leading to this consensus answer are assigned a positive reward ($r=1$), while others receive zero. Crucially, this reward ignores the semantic diversity or logical validity of the intermediate reasoning steps (Rationale), making it a coarse signal that relies heavily on the diversity of the generated candidates to be effective. All parameter updates are applied via LoRA for computational efficiency.
Results.
Figure 5 illustrates the reward curves evaluated every 5 steps. We compare the standard Qwen3-1.7B model using high-temperature sampling (temp=1.0) against B-Trans. The B-Trans exhibits high variance and oscillation but achieves significantly higher peak accuracy compared to the baseline.
Figure 5: Accuracy curve during training. Reported at every five training steps.
Analysis.
Two key insights emerge from this experiment. First, the hyperparameter $\sigma$ serves as an effective control for exploration intensity. Higher values allow the model to escape local optima through aggressive exploration, akin to brainstorming wild ideas, which is necessary for the challenging AIME24 problems. Second, the success of B-Trans in the TTRL setting suggests an emergent self-correction capability. This parallels human cognitive processes, where internal deliberation and re-evaluation can refine understanding even without external feedback. Even with a coarse consensus-based reward, B-Trans effectively filters out homogeneous cognition common to single-model instances by providing a broader set of rationales for the vote.
6 Ablation Study and Discussion
In this section, we dissect the architectural choices of B-Trans to validate our core design principles. We systematically investigate the necessity of temporal consistency in noise injection, justify the decision to perturb biases over weights, and verify the robustness of our approach across different post-training frameworks.
6.1 Temporal Consistency: Sequence-wise vs. Token-wise Noise
@@ 6 Ablation Study and Discussion
To validate the necessity of hypothesis sampling (Section 3.2), we compare B-Trans against a Token-wise Noise baseline on MATH-500 using Qwen3-1.8B. While B-Trans samples a single latent $z$ per sequence, the baseline resamples $z_{t}\sim\mathcal{N}(0,\sigma^{2}I)$ at every forward pass. Crucially, we fix an identical noise magnitude ($\sigma=0.02$) for both settings to isolate the impact of temporal inconsistency.
@@
Results & Analysis. The Token-wise baseline suffers catastrophic degradation. Despite identical noise levels, its SCS drops to 0.42 (vs. 0.58 for B-Trans), correlating with a decrease in accuracy (40.2% vs. 46.4%). This empirical failure corroborates the theoretical insight established by gal2016dropout at RNNs/LSTMs age: valid variational inference requires the stochastic mask to remain invariant across time steps to represent a single coherent hypothesis. B-Trans satisfies this “Temporal Consistency” requisite through lightweight bias perturbation, effectively capturing the benefits of weight uncertainty without the memory overhead of full ensemble masks.
6.2 Post-training Infrastructure Agnosticism
To validate that B-Trans serves as a "plug-and-play" module—agnostic not just to the model architecture but also to the training infrastructure, we repeat our TRL-based vonwerra2022trl RLVR experiments (Section 5.2) using another widely adopted RL engine, namely VeRL sheng2024hybridflow (Volcano Engine Reinforcement Learning). We observed no statistically significant difference in the final convergence performance or learning dynamics between the two implementations. While the VeRL implementation exhibited faster wall-clock training speeds in our environment, we attribute this to differences in rollout acceleration libraries rather than the B-Trans mechanism itself. We expect B-Trans to be agnostic to the underlying base LLMs and training infrastructure, supporting its potential as a surgical plug-and-play approach for diverse use cases.
7 Future Work
Our current implementation models the prior distribution as a simple isotropic Gaussian with a scalar hyperparameter $\sigma$. While this serves as an effective proof-of-concept, the assumption of uniform uncertainty across all layers and dimensions is a strong simplification.
A natural progression for this line of research is to develop a learnable or hierarchical prior. In this initial exploration, we adhered to a fixed prior to mitigate the risks associated with post-training optimization. Learning variance parameters on limited downstream datasets can lead to overfitting, where the induced population loses its general competence. We anticipate that future efforts, particularly those with sufficient computational resources to train on large-scale general corpora, could successfully learn these uncertainty parameters. Such a data-driven prior would likely act as a more robust “population,” capturing layer-specific sensitivities that a scalar $\sigma$ cannot.
@@ 8 Conclusion
We revisit the prevailing view of Large Language Models as static point estimates and argue that lightweight parameter-space uncertainty can be practically useful for exploration and diversity. We introduce B-Trans, a population-based framework that samples diverse yet coherent model instances from a single Transformer backbone using a variationally motivated Gaussian posterior proxy. By applying temporally consistent perturbations to normalization biases, B-Trans simulates the wisdom of crowds at negligible cost, yielding significant gains in zero-shot diversity and sparse-reward reinforcement learning. Our results suggest that representing uncertainty as a sampleable local proxy, even in a simple form, can enable population-driven reasoning within a single pre-trained model.
@@
While B-Trans demonstrates a promising direction for inducing implicit ensembles, our current study is intentionally scoped as a minimal, representative proof-of-concept centered on architectural simplicity and controlled comparisons. We do not perform extensive hyperparameter sweeps, broad architecture coverage, or large scale benchmarking beyond validating the key design principles. Our intervention is restricted to the bias terms of normalization layers, which is a simplified proxy rather than a full Bayesian neural network. While this surgical approach is computationally negligible, it is a simplified approximation of a full Bayesian Neural Network. The weights in the attention and feed-forward layers remain deterministic, potentially limiting the expressivity of the induced posterior.

### v2

@@ 4.3 Temporal Consistency via Noise Caching
As mentioned in Section 3.2, for each sequence generation, we sample one LLM (analogous to a distinct individual from the population) to solve the problem. Thus, we implement a sequence-level noise caching mechanism (see Figure 1). As detailed in the BayesianBiasWrapper, the process ensures that the stochastic perturbation defines a stable reasoning flow for the entire response: (1) Initialization: Upon receiving the prompt, we sample the latent variable $\mathbf{z}$ once and cache it, effectively drawing a single sample $\theta^{(k)}$ from the approximate posterior. (2) Generation: For all subsequent autoregressive steps within the same sequence, the model applies this frozen $\mathbf{z}$ to the normalization layers.
This design ensures that the stochasticity is strictly inter-sequence rather than intra-sequence. Crucially, unlike full weight-masking methods (e.g., MC Dropout), our implementation is computationally negligible, requiring only an element-wise addition of a cached tensor without the overhead of generating large binary masks. Specifically, implementing theoretically correct (frozen) MC Dropout on a 7B model would require caching around 14GB of masks per user instance, assuming pre-expanding the binary masks to FP16 to align with weight precision to avoid prohibitive on-the-fly decompression overheads. In contrast, B-Trans only caches the bias noise vectors (<1MB), reducing the memory footprint by orders of magnitude while maintaining temporal consistency.
@@ 5 Experiments
To validate that B-Trans can produce diverse but meaningful minds from one model, we conduct experiments under two settings: (1) B-Trans for zero-shot generation, focusing on improving semantic diversity and also pass@k for accuracy, and (2) B-Trans for output sampling, focusing on providing a useful learning signal for Reinforcement Learning with Verifiable Rewards (RLVR). Across both settings, our goal is to provide a minimal proof-of-concept for population-based exploration under controlled, standard protocols, rather than an exhaustive benchmark.
5.1 Zero-Shot Diversity and Creativity
The premise of B-Trans is that each sampled response corresponds to a different model instance. We evaluate whether B-Trans yields meaningful heterogeneity compared to the deterministic baseline with high-temperature decoding.
@@ 5.1 Zero-Shot Diversity and Creativity
Figure 4: PCA visualization of response embeddings from Qwen3 and Llama-3.1 models. The query is “Output a hard question to humanity (super concise and short), independent of theme.” Standard sampling confines models to small semantic regions. B-Trans significantly expands the territories, allowing the model to escape its default mode and even explore semantic regions typically associated with the other models (highlighted in red box).
@@
We study intrinsic population behavior on two benchmarks with different measurement goals. On MMLU-Pro Wang et al. (2024) (a challenging scientific reasoning dataset that contains 12K complex questions across various disciplines), we use Pass@K to test whether additional sampled individuals provide valid reasoning paths: for each question, we sample $K$ independent LLM instances and count the answer as correct if any sample is correct. Thus, improved scaling with larger $K$ indicates that the induced diversity is not mere repetition but yields a correct solution. On INFINITY-CHAT Jiang et al. (2025) (100 open-ended questions that admit a wide range of plausible answers with no ground-truth correctness), we measure semantic diversity using average pairwise embedding cosine distance Reimers and Gurevych (2019); Song et al. (2020) and visualize the embedding geometry with PCA following Artificial Hivemind analysis (Jiang et al., 2025). The goal is to quantify and visualize semantic heterogeneity. For both tasks, we contrast B-Trans against two distinct base models, Qwen3 1.7B, 8B Yang et al. (2025), and Llama-3.1 Grattafiori et al. (2024). All models share the same decoding strategies.
@@ Results.
On MMLU-Pro (Figure 2), while B-Trans matches the baseline at $K=1$, it demonstrates significantly stronger scaling behavior as the sample budget increases. This trend confirms that the diversity introduced by B-Trans is semantically meaningful: rather than generating redundant repetition (which would yield diminishing returns), the B-Trans population successfully explores valid reasoning trajectories that are not found by the deterministic baselines. Complementing this, the embedding distance analysis on INFINITY-CHAT (Figure 3) further quantifies this behavior, confirming better output heterogeneity.
Qualitative Analysis
Figure 4 visualizes the semantic landscape, where spatial proximity indicates the semantic similarity of response embeddings (dots). Note that the horizontal distance reflects the most fundamental conceptual differences between each response.
Under standard sampling (temp-0.5), Qwen3 and Llama-3.1 occupy disjoint, narrow regions in the semantic space, reflecting their distinct pre-trained biases. While high temperature (temp-1) slightly expands these clusters, it fails to alter their fundamental topology. In contrast, B-Trans generates a significantly broader distribution, forming distinct clusters that explore diverse reasoning paths.
Notably, B-Trans can occasionally reach semantic regions that are atypical for the original base model. As examples highlighted in the red box, Qwen3-8B-btrans generates one response that is very close to Llama-3.1’s style. This validates our goal of mining different minds from one model.
@@ 5.2 Exploration in Reinforcement Learning with Verifiable Rewards (RLVR)
In Reinforcement Learning, the critical bottleneck is often the exploration-exploitation trade-off. Standard LLM policies rely on action-space noise (sampling tokens) to explore. In this section, we show that parameter-space noise (sampling models) provides a more consistent and deeper form of exploration under the setting of RLVR.
Setting.
We follow the default configuration of the SimpleRL-Zoo framework as-is, which is a popular RLVR recipe to improve models’ reasoning abilities using Group Relative Policy Optimization (GRPO) Guo et al. (2025). Our studies include different sizes of Qwen3 models (1.7B and 8B), math-specialized Qwen2.5-Math-7B, and Llama3.1-8B. For each model, we compare two approaches to sample a group of responses (16 outputs) for each input for GRPO reward computing : (1) using a deterministic base model, and (2) using B-Trans. Note that we keep the policy model the same for both settings and focus solely on the difference in sampling strategies. To save computing, we only update Low-Rank Adaptation (LoRA) Hu et al. (2022) adapters for the policy model instead of full-parameter fine-tuning (less than 1% trainable parameters), and we train on 400 examples sampled from the same distribution of the SimpleRL-Zoo dataset via difficulty-stratified sampling. We test the finetuned model on three standard mathematical reasoning datasets, i.e., GSM8K Lewkowycz et al. (2022), MATH-500 Lightman et al. (2023), and Minerva Math Lewkowycz et al. (2022), where the ground truth is verifiable.
Results.
As shown in Table 1, B-Trans achieves consistent gains across all four benchmarks. GRPO relies on informative relative comparisons within the sampled group. When the rollout group is homogeneous, relative advantages become less informative, and updates can stagnate. By producing group members from distinct parameter-space instances, B-Trans increases the chance that at least one candidate reaches a higher-reward region, strengthening the learning signal.
@@ Results.
Table 1: Accuracy (%) on RLVR tasks using GRPO. B-Trans consistently outperforms the baseline.
6 Ablation Study
@@
To validate the necessity of hypothesis sampling (Section 3.2), we compare B-Trans against a Token-wise Noise baseline on MATH-500 using Qwen3-1.7B. While B-Trans samples a single latent $z$ per sequence, the baseline resamples $z_{t}\sim\mathcal{N}(0,\sigma^{2}I)$ at every forward pass. We fix an identical noise magnitude ($\sigma=0.02$) for both settings to isolate the impact of temporal inconsistency.
@@
Results and Analysis
The Token-wise baseline suffers catastrophic degradation. Despite identical noise levels, its SCS drops to 0.42 (vs. 0.58 for B-Trans), correlating with a decrease in accuracy (40.2% vs. 46.4%). This empirical failure corroborates the theoretical insight established by Gal and Ghahramani (2016) at RNNs/LSTMs age: valid variational inference requires the stochastic mask to remain invariant across time steps to represent a single coherent hypothesis. B-Trans satisfies this Temporal Consistency requisite through lightweight bias perturbation, effectively capturing the benefits of weight uncertainty without the memory overhead of full ensemble masks.
7 Future Work and Discussion
Our current implementation models the prior distribution as a simple isotropic Gaussian with a scalar hyperparameter $\sigma$. While this serves as an effective proof-of-concept, the assumption of uniform uncertainty across all layers and dimensions is a strong simplification. Furthermore, as mentioned in Section 4.2, we could also introduce a distribution over parameters other than Bias-like offset, and the weights in normalization could be a natural target.
Another progression for this line of research is to develop a learnable or hierarchical prior. In this initial exploration, we adhered to a fixed $\sigma$ to mitigate the risks associated with post-training optimization. Learning variance parameters on limited downstream datasets can lead to overfitting, where the induced population loses its general competence. Ideally, the population model can be learned from a large corpus reflecting real population behavior, similar to using real human interaction data for Bayesian recommender system training Zhang and Koren (2007). We anticipate that future efforts, particularly those with sufficient computational resources to train on large-scale general corpora, could successfully learn these uncertainty parameters. Such a data-driven prior would likely act as a more robust population, capturing layer-specific sensitivities that a scalar $\sigma$ cannot.
@@ 8 Conclusion
We revisit the prevailing view of Large Language Models as static point estimates and argue that lightweight parameter-space modification is a practical catalyst for exploration and diversity. Benefited from the wisdom of crowds, our experiments demonstrate that B-Trans, even with a simple instantiation, provides meaningful and diverse responses at negligible cost. In an era where foundation models have achieved strong deterministic prediction capabilities (exploitation), we envision that B-Trans contributes to the next generation of models by breaking through structural homogenization and facilitating efficient learning via deeper exploration.
@@
While B-Trans demonstrates a promising direction for inducing implicit ensembles, our current study is intentionally scoped as a minimal, representative proof-of-concept centered on architectural simplicity and controlled comparisons. We do not perform extensive hyperparameter sweeps, broad architecture coverage, or large-scale benchmarking beyond validating the key design principles. Our intervention is restricted to the biased terms of normalization layers, which is a simplified proxy rather than a full Bayesian neural network. While this surgical approach is computationally negligible, it is a simplified approximation of a full Bayesian Neural Network. The weights in the attention and feed-forward layers remain deterministic, potentially limiting the expressivity of the induced posterior.
Ethics Considerations and Broader Impact
This work introduces a parameter-space sampling method to enhance LLM diversity during inference. Regarding potential risks, we acknowledge that techniques improving generative diversity could theoretically be misused to generate varied disinformation or spam. However, our primary objective is to enable deeper exploration in reasoning tasks. Environmentally, B-Trans offers a positive impact by approximating population-level intelligence from a single model, thereby avoiding the significant carbon footprint associated with training and maintaining full model ensembles.